OpenStudy (anonymous):

Are all these numbers composite: 961, 1961, 2961, 3961, 4961, 5961, 6961, 7961, 8961, 9961

OpenStudy (anonymous):

So what do you know about prime and composite numbers?

OpenStudy (anonymous):

if it can go into more than 1 and itself evenly, it is prime; if not, it is composite. if you try to factor the number, 961, does anything go into it? 961: 1, 961-composite 1961: 1, 1961- composite 2961: 1, 2961- composite 3961: 1, 3961- composite 4961: 1, 4961- composite 5961: 1, 5961- composite but 6961 is prime.

OpenStudy (wolf1728):

A composite number is one that is a multiple of at least two numbers other than itself and 1. There are no general laws regarding testing a prime and no formulas that will continuously generate primes. Generally, you can test if a number ISN'T prime if • it is even - applies to all numbers except 2 • it is a multiple of 5 except of course 5. So, with those 2 rules we can definitely say that 60% of all numbers are NOT prime (even numbers, and numbers ending in 5). • If the sum of all the digits (example 2,961) reduces to 3 , 6 or 9 it is NOT a prime. 2+9+6+1 = 18 = 1+8=9 not a prime. For a list such as yours the easiest thing to do would be to check those numbers against a list of prime numbers. Also, you could go to an online factoring program to test the number. You could test a number by dividing it by all primes up to the square root of a number but that is time consuming. For example, 7,961 is 19 * 419 and even that would take a little while to test.

OpenStudy (anonymous):

"• If the sum of all the digits (example 2,961) reduces to 3 , 6 or 9 it is NOT a prime. 2+9+6+1 = 18 = 1+8=9 not a prime" didnt know this b4 wolf, cheers!